OpenStudy (ucreategames):

35x^-2 + 12x^-1 + 1 = 0... Solve for x

OpenStudy (dangerousjesse):

There are two possible answers when you work it out.. \(\large\frac{35}{x^2}+\frac{12}{x}+1 = 0\)

OpenStudy (dangerousjesse):

I keep misreading the equation, sorry.

OpenStudy (ucreategames):

Lol its okay...

OpenStudy (ucreategames):

You're doing great. I posted this problem to see who could get the answer. It's not for me.

OpenStudy (dangerousjesse):

Rewrite the left hand side of the equation. \(1+\frac{12}{x}+\frac{35}{x^2} = 1+\frac{12}{x}+\frac{35}{x^2:}\) \(1+\frac{12}{x}+\frac{35}{x^2} = 0\) Write the left hand side as a single fraction. Bring \(\frac{1+12}{x+35}\div x^2\) together using the common denominator x^2: \(\frac{x^2+12 x+35}{x^2} = 0\) Multiply both sides by a polynomial to clear fractions. Multiply both sides by \(x^2:\) \(x^2+12 x+35 = 0\) Factor the left hand side. The left hand side factors into a product with two terms: \((x+5) (x+7) = 0\) Solve each term in the product separately. Split into two equations: \(x+5 = 0\) or \(x+7 = 0\) Look at the first equation: Solve for x. Subtract 5 from both sides: \(x = -5\) or \(x+7 = 0\) Look at the second equation: Solve for x. Subtract 7 from both sides...

OpenStudy (dangerousjesse):

\(x=-7\) or \(-5\)

OpenStudy (ucreategames):

You're correct...!